OpenStudy (cenaida):

Solve the following system of equations using elimination: 6x - 5y = 3 4x - 36y = 100 A. (-3, -2) B. (-2, -3) C. (-2, 3) D. (2, 3)

OpenStudy (cenaida):

@abb0t

OpenStudy (abb0t):

Ok, let's start by naming them equation 1, and equation 2. yes?

OpenStudy (cenaida):

yes

OpenStudy (abb0t):

The goal here is to rearrage one of them so you can substitute into the other. If you rearrange eqation 1 to substitue say...into \(x\). you can proceed like this: x = \(\sf \color{red}{\frac{1}{2}+\frac{5}{6}y}\)

OpenStudy (abb0t):

Now, notice how you have x = some function. Yes?

OpenStudy (cenaida):

yess I do

OpenStudy (abb0t):

Substitute THAT function in \(\sf \color{red}{red}\) into \(x\) on equation 2! You will notice that you have a function in terms of oNLY y! So you have: 4(\(\sf \color{red}{\frac{5}{6}y+\frac{1}{2}}\))-36y=100

OpenStudy (abb0t):

Now, distribute the \(\sf \color{red}{4}\) into the parenthesis and solve for \(\sf \color{blue}{y}\)!! Then when you have a value for y, you can plug in that value into equation 1, and solve for \(x\)! and you're done. Then you have two values for x, and y!. You can check if they are correct by plugging those values into ANY of the two equations. You should get 3, OR 100 if you plug those values in for x, or y.

OpenStudy (anonymous):

@abb0t , you better get this one right

OpenStudy (anonymous):

by the way, nice hair

OpenStudy (abb0t):

you should come over and comb it sometime @Luis_Rivera

OpenStudy (anonymous):

lol, no thanks, let nincomb do it

OpenStudy (cenaida):

hahaha so is I D?

OpenStudy (cenaida):

is it? @Luis_Rivera @abb0t

OpenStudy (cenaida):

2

OpenStudy (abb0t):

Does: 6(\(\sf\color{red}{2}\))-5(\(\sf\color{red}{3}\))=3???

OpenStudy (cenaida):

-3

OpenStudy (abb0t):

Yes. And does \(\sf\color{red}{-3}\) = 3?

OpenStudy (anonymous):

get to the point!

OpenStudy (cenaida):

no?..

OpenStudy (abb0t):

Your answer is B!

OpenStudy (cenaida):

oh yeaa! thanks!

OpenStudy (anonymous):

yes @cenaida , the girl with the red hair is correct

OpenStudy (cenaida):

hahaa lol